OpenStudy (anonymous):

A positive integer is divided by a smaller positive integer. The integer part of the quotient is 7 and the remainder is 2. If the sum of the two integers is 42, what are the two integers?

OpenStudy (shubhamsrg):

let greater one be x and smaller be y it says that x = 7y + 2 and x+y =42.. shouldnt it be easy now ?

OpenStudy (anonymous):

can you put it into an equation? @shubhamsrg

OpenStudy (shubhamsrg):

what ? you mean one single eqn ? just make this subsn : put the value of x from 1st eqn into 2nd one..and see the magic! :D

OpenStudy (anonymous):

@shubhamsrg ,only the integer part of the quotient is given........ no problem kyaaa!!!

OpenStudy (anonymous):

i dont understand

OpenStudy (shubhamsrg):

well if you take that into account then i dont think you can actually come up with 2 eqns.. how will you use this fact..and at this level,,i guess we have to take that as 7 only..hmmn,.

OpenStudy (shubhamsrg):

@tbrooks3 see we have this x=7y+2 and x+y =42 from 1st eqn,,2nd becomes => (7y+2) +y =42 will this help ?

OpenStudy (anonymous):

yes thanks

OpenStudy (anonymous):

i too thought so... i asked just for clarificaton:)))

OpenStudy (shubhamsrg):

glad to help ^_^

OpenStudy (shubhamsrg):

@anusha.p lets consider what you're saying,,let the fraction part be f now we have x = (7+f)y + 2 and x+y =42 => 8y + fy = 40 y = 40/(8 +f) since f lies between 0 and 1,,y thus lies between 5 and 4.444... thus x lies between 37 and 38.5555... so you see..hmmn..

OpenStudy (anonymous):

@shubhamsrg got u!!!